OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (chriss):

x+x+1+x+2=267 3x+3=267 3x=264 x=88 x+1=89 x+2=90 88+89+90=267

OpenStudy (anonymous):

?

OpenStudy (chriss):

sorry, didn't catch the wording at the end of the question... the largest integer is 90.

OpenStudy (anonymous):

Thanks

OpenStudy (chriss):

yw